### v1

@@ 0.3.4 Training
The resulting customised PGD training seeks to identify the worst perturbations within the custom-defined subspace, and trains the given neural network to classify those perturbations correctly, in order to make the network robust to adversarial inputs in the chosen subspace.
0.3.5 Choice of Verification Algorithm
As stated earlier, our approach in this study involves the utilization of cutting-edge tools for DNN verification. Initially, we employ ERAN [110], a state-of-the-art abstract interpretation-based method. This choice is made over IBP due to its ability to yield tighter bounds. Subsequently, we conduct comparisons and integrate Marabou [131], a state-of-the-art complete and precise SMT-based verifier. This enables us to attain the highest verification percentage, maximizing the tightness of the bounds.
We will use notation $\mathcal{V}$ to refer to a verifier abstractly.
0.4 Characterisation of Verifiable Subspaces
  • We start with introducing the metric of generalisability of (verified) subspaces, and introducing the problem of the generalisability-verifiability trade-off.
  • We show that the use of semantic subspaces helps to find a better balance between generalisability and verifiability, as compared to the use of geometric subspaces.
0.4.1 Generalisability-Verifiability Trade-off
This subsection defines the new metric of generalisability and shows its effect on (naively defined) geometric subspaces.
Generalisability of Verified Subspaces: Formal Definition
Let us start with recalling the existing standard metrics used in DNN verification. Recall that we are given an NLP dataset $\mathcal{Y}=\{s_{1},\ldots,s_{q}\}$, moreover we assume that each $s_{i}$ is assigned a correct class from $C=\{c_{1},\ldots,c_{n}\}$. We restrict to the case of binary classification in this paper for simplicity, so we will assume $C=\{c_{1},c_{2}\}$. Furthermore, we are given an embedding function $\mathbb{E}:\mathcal{Y}\rightarrow^{m}$, and a DNN $N:^{m}\rightarrow^{n}$. Usually $n$ corresponds to the number of classes, and thus in case of binary classification, we have $N:^{m}\rightarrow^{2}$. Classification, or assignment of a vector $v$ in ${}^{m}$ to a class, depends on the highest value in the vector $N(v)$. The most popular metric is accuracy of $N$, which is measured as a percentage of $q$ vectors in ${}^{m}$ that are assigned to a correct class by $N$. Note that this metric checks a finite number of points in ${}^{m}$ given by the data set.
The most popular metric in DNN verification is verifiability. Recall that we can define subspaces $\{\mathcal{S}_{1},\ldots,\mathcal{S}_{s}\}$ of ${}^{m}$; in such a way that each of them is associated with a class (to which all points in that subspace should belong). A DNN verifier $\mathcal{V}$ takes an $\mathcal{S}_{i}$, its designated class $c_{i}$ and $N$ as an input, and returns as an output an answer whether all points in $\mathcal{S}_{i}$ are guaranteed to be assigned to the class $c_{i}$ by $N$. Verifiability is a percentage of subspaces in $\{\mathcal{S}_{1},\ldots,\mathcal{S}_{s}\}$ that are successfully verified. All DNN verification papers report this measure. Note that each $\mathcal{S}_{i}$ contains an infinite number of points.
We are introducing a third metric – generalisability of (verified) subspaces. Suppose we have a subspace $\mathcal{S}_{i}$ that verifiably consists only of vectors that are assigned to a class $c_{i}$ by $N$. Because of the embedding gap, we cannot know in advance how many valid sentences in $\mathcal{Y}$ (or outside of $\mathcal{Y}$ !) will be mapped into $\mathcal{S}_{i}$ by $\mathbb{E}$. Checking the former is rather easy, but may not reflect the idea of generalisability to unseen similar sentences. Checking the latter is hard. We propose the following effective heuristic. By the constructions of the previous section, any $\mathcal{S}_{i}$ necessarily contains an embedding of at least one element $s_{i}$ from $\mathcal{Y}$. Choosing any perturbation algorithm $\mathcal{P}$ (of kind $k$), we can form a set $\mathcal{A}^{k}_{s_{i}}$, as described in Section 0.3.3. Note that $\mathcal{P}$ (and $k$) can be given by a collection of different perturbation algorithms and their kinds. The key assumption is that $\mathcal{A}^{k}_{s_{i}}$ contains valid sentences semantically similar to $s_{i}$ and belonging to the same class.
Consider a set $\mathcal{V}_{i}$ of vectors in ${}^{m}$ such that each vector $v_{j}\in\mathcal{V}_{i}$ is defined as $v_{j}=\mathbb{E}(s_{j})$, with $s_{j}\in\mathcal{A}^{k}_{s_{i}}$. This is the set of embeddings of sentences contained in $\mathcal{A}^{k}_{s_{i}}$. Percentage of elements of $\mathcal{V}_{i}$ contained in $\mathcal{S}_{i}$ gives us the generalisability of $\mathcal{S}_{i}$. Note that each $\mathcal{V}_{i}$ is finite, because each $\mathcal{A}^{k}_{s_{i}}$ is finite.
Note that, unlike accuracy and verifiability, the generalisability metric does not explicitly depend on any DNN. However, in this paper we only study generalisability of verifiable subspaces, and thus the existence of a verified DNN $N$ will be assumed.
Because verifiability is reported as a percentage of $\{\mathcal{S}_{1},\ldots,\mathcal{S}_{s}\}$, it will be convenient to report generalisability over the set of subspaces $\mathcal{S}=\bigcup\mathcal{S}_{i}$. For this, we take $\mathcal{A}=\bigcup\mathcal{A}^{k}_{s_{i}}$ and the corresponding $\mathcal{V}=\bigcup\mathcal{V}_{i}$ for all $s_{i}\in\mathcal{Y}$ and compute generalisability of $\mathcal{S}$ as a percentage of elements of $\mathcal{V}$ contained in all subspaces of $\mathcal{S}$ [^6] Note that this calculation allows an element of $\mathcal{A}^{k}_{s_{i}}$ to count towards generalisability if it belongs to $\mathcal{S}_{j}$, regardless of whether $i=j$ or $i\neq j$. Different approaches might involve restraining the validity to $i=j$, or computing generalisability of each $\mathcal{S}_{j}$ individually and taking an average. The current choice favours a global view by calculating how generalisable the whole collection $\mathcal{S}$ is. . When it is important to emphasise the kind of perturbation, we will use the notation $\mathcal{A}_{k}$ for $\bigcup\mathcal{A}^{k}_{s_{i}}$.
Base Line Experiments: Understanding the Properties of Embedding Spaces
Figure 3: An example of hyper-rectangle drawn around all points of the same class (left), shrunk hyper-rectangle $\mathbb{H}_{sh}$ that is obtained by excluding all points from the opposite class (center) and clustered hyper-rectangles (right) in 2-dimensions. The red dots represent sentences in the embedding space of one class, while the blue dots are embedded sentences that do not belong to that class.
The methodology defined thus far has given basic intuitions about the relative nature of the NLP verification pipeline. Bearing this in mind, it is important to start our analysis with the general study of the embedding (sub)spaces and suitable base line settings.
Benchmark data sets will be abbreviated as “RUAR" and “Medical". For a benchmark network $N:^{m}\rightarrow^{2}$, we train a medium-sized fully-connected DNN, using stochastic gradient descent and cross-entropy loss. The main requirement for a benchmark network is its sufficient accuracy, see Table 5.
$N_{base}$
No
0.9387 ± 0.0014
0.9357 ± 0.0018
0.9632 ± 0.0005
0.9449 ± 0.0026
Table 5: Accuracy of the baseline DNN on the R-U-A-Robot and the Medical datasets.
For the choice of subspaces, which is our main interest in this paper, two extreme (and trivially defined) benchmarks, are the following geometric subspaces:
  • a subspace drawn around all embedded sentences of the same class in $\mathcal{Y}$. This is the largest subspace one may wish to verify but we should assume that verifiability of such a subspace would be near $0\%$. It is illustrated in the first graph of Figure 3.
@@ 0.4.1 Generalisability-Verifiability Trade-off
We need to first understand exact geometric properties (e.g. volume, $\epsilon$ values) and exact verifiability figures for these two extremes. Let us start with understanding the volumes.
Volumes of Subspaces.
To start with the largest hyper-rectangle, we need to ensure that it is shrunk to exclude all samples from $\mathcal{Y}$ from the wrong class. We define the shrinking method that involves reducing the size of the hyper-rectangle to eliminate each input within the hyper-rectangle that does not pertain to the chosen class. Second graph of Figure 3 gives a visual intuition of how this is done.
Formally, suppose we have a hyper-rectangle $\mathbb{H}_{\mathcal{X}}$ and a vector $\hat{x}\in\mathbb{H}_{\mathcal{X}}$ that violates the class requirement. Recall that by definition:
$\mathbb{H}_{\mathcal{X}}\coloneqq\{x\in^{m}:\forall j\in[1\;,\;...\;,\;m]\;.\;\>x^{j}\in[\min\nolimits_{i}(\mathcal{X}^{ij}),\;\;\max\nolimits_{i}(\mathcal{X}^{ij})]\}.$
The algorithm starts with choosing a dimension $j\in[1,...,m]$, and computing whether $|\hat{x}^{j}-\min\nolimits_{i}(\mathcal{X}^{ij})|\leq|\hat{x}^{j}-\max\nolimits_{i}(\mathcal{X}^{ij})|$. If the inequality holds, take $d_{\min\nolimits}=|\hat{x}^{j}-\min\nolimits_{j}(\mathcal{X}^{ij})|$ and compute a new minimum for the dimension $j$: $\min\nolimits^{new}_{i}(\mathcal{X}^{ij})=\min\nolimits_{i}(\mathcal{X}^{ij})+d_{min}+\delta$, where $\delta$ is a small positive number (we use $e^{-100}$). We then recompute $\mathbb{H}_{sh}=\mathbb{H}_{\mathcal{X}}$ with $\min\nolimits_{i}(\mathcal{X}^{ij})=\min\nolimits^{new}_{i}(\mathcal{X}^{ij})$ for the chosen dimension $j$.
If the inequality does not hold, take $d_{\max\nolimits}=|\hat{x}^{j}-\max\nolimits_{i}(\mathcal{X}^{ij})|$ and compute a new maximum for the dimension $j$: $\max\nolimits^{new}_{i}(\mathcal{X}^{ij})=\max\nolimits_{i}(\mathcal{X}^{ij})-d_{max}-\delta$, with $\delta$ defined as above. We then recompute $\mathbb{H}_{sh}=\mathbb{H}_{\mathcal{X}}$ with $\max\nolimits_{i}(\mathcal{X}^{ij})=\max\nolimits^{new}_{i}(\mathcal{X}^{ij})$ for the chosen dimension $j$.
Note that modifying just one dimension of $\hat{x}$ like this guarantees that it is not in $\mathbb{H}_{sh}$. The algorithm we apply when computing $\mathbb{H}_{sh}$ involves a heuristic for choosing the optimal dimension along which we modify the hyper-rectangle. The idea is to make the most conservative choice, that excludes as few other embedded sentences as possible [^7] Note that this algorithm shrinks exactly one dimension by a minimal amount to exclude the unwanted embedded sentence. This choice keeps the algorithm fast while guaranteeing the subspace to retain the highest number of wanted inputs. However, it is not necessarily the best choice for verification: there might be cases where perturbations of the unwanted input are left inside after shrinking and, if the network classifies them correctly, the subspace can never be verified. For large subspaces, our algorithm might render verification unachievable and more clever algorithms should be explored and discussed.. It amounts to computing how many inputs from the chosen class remain inside of the hyper-rectangle after applying the shrinking algorithm to one $j\in[1,...,m]$ at a time, separately. We choose the dimension $j$ for which the hyper-rectangle retains the most inputs of the correct class. The algorithm is run iteratively for every element of $\mathbb{H}_{sh}$ that has the wrong class, until $\mathbb{H}_{sh}$ contains only embeddings of sentences of one class.

### v3

@@ 4.3 Verifiability-Generalisability Trade-off for Geometric Subspaces
Table 8: Sets of geometric subspaces used in the experiments, their cardinality and average volumes of hyper-rectangles. All shapes are eigenspace rotated for better precision.
4.3.1 Verifiability of Geometric Subspaces
Next, we pass each set of hyper-rectangles and the given network to the ERAN verifier and measure verifiability. Table 9 shows that, as expected, the shrunk hyper-rectangle $\mathbb{H}_{sh}$ achieves 0% verifiability, and the various clustered hyper-rectangles ($\mathbb{H}_{50}$, $\mathbb{H}_{100}$ $\mathbb{H}_{200}$, $\mathbb{H}_{250}$) achieve at most negligible verifiability. In contrast, the baseline $\mathbb{H}_{\epsilon=0.005}$ achieves up to $99.60\%$ verifiability. This suggests that $\epsilon=0.005$ is a good benchmark for a different extreme. Table 8 can give us an intuition of why $\mathbb{H}_{\epsilon=0.005}$ has notably higher verifiability than the other hyper-rectangles: the volume of $\mathbb{H}_{\epsilon=0.005}$ is several orders of magnitude smaller. We call this effect low verifiability of the high-volume subspaces.
Dataset Model $\boldsymbol{\mathbb{H}_{sh}}$ $\boldsymbol{\mathbb{H}_{50}}$ $\boldsymbol{\mathbb{H}_{100}}$ $\boldsymbol{\mathbb{H}_{200}}$ $\boldsymbol{\mathbb{H}_{250}}$ $\boldsymbol{\mathbb{H}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}_{\epsilon=0.005}}$
RUAR $N_{base}$ 0.00% 0.00% 1.33% 0.52% 0.41% 0.00% 88.67%
Medical $N_{base}$ 0.00% 0.00% 0.00% 2.10% 4.08% 5.00% 97.86%
Table 9: Verifiability of the baseline DNN on the RUAR and the Medical datasets, for a selection of geometric subspaces; using the ERAN verifier.
Tables 8 and 9 suggest that smaller subspaces are more verifiable. One may also conjecture that they are less generalisable (as they will contain fewer embedded sentences). We now will confirm this via experiments; we are particularly interested in understanding how quickly generalisability deteriorates as verifiability increases.
4.3.2 Generalisability of Geometric Subspaces
To test generalisability, we algorithmically generate a new dataset $\mathcal{A}^{b}_{t}(\mathcal{Y}^{pos})$ containing its semantic perturbations, using the method described in Section 3.1. The choice to use only positive sentences is motivated by the nature of the chosen datasets - both Medical and RUAR sentences split into:
  • a negative class that represents “all other sentences”. These “other sentences” are not grouped by any specific semantic meaning and therefore do not form one coherent semantic category.
However Section 5 will make use of $\mathcal{A}^{b}_{t}(\mathcal{Y}^{neg})$ in the context of discussing the embedding error of verified subspaces.
@@ 4.3.2 Generalisability of Geometric Subspaces
Dataset Experiment Avg. Volume of hyper-rectangles Generalisability (%) Number of sentences contributing to generalisability Total Sentences in $\mathcal{A}^{b}_{t}(\mathcal{Y}^{pos})$
RUAR $\mathbb{H}_{\epsilon=0.005}$ 1.00e-60 1.95 2821 144500
$\mathbb{H}_{\epsilon=0.05}$ 1.00e-30 38.47 55592 144500
$\mathbb{H}_{sh}$ 7.55e-11 50.91 73561 144500
Medical $\mathbb{H}_{\epsilon=0.005}$ 1.00e-60 0.09 10 11209
$\mathbb{H}_{\epsilon=0.05}$ 1.00e-30 28.49 3194 11209
$\mathbb{H}_{sh}$ 2.6e-09 37.13 4162 11209
Table 10: Generalisability of the selected geometric subspaces $\mathbb{H}_{\epsilon=0.005}$, $\mathbb{H}_{\epsilon=0.05}$ and $\mathbb{H}_{sh}$, measured on the sets of semantic perturbations $\mathcal{A}^{b}_{t_{RAUR}}(\mathcal{Y}^{pos})$ and $\mathcal{A}^{b}_{t_{medical}}(\mathcal{Y}^{pos})$.
Table 10 shows that the most verifiable subspace $\mathbb{H}_{\epsilon=0.005}$ is the least generalisable. This means $\mathbb{H}_{\epsilon=0.005}$ may not contain any valid new sentences apart from the one for which it was formed! At the same time, $\mathbb{H}_{\epsilon=0.05}$ has up to $48\%$ of generalisability at the expense of only up to $5\%$ of verifiability (cf. Table 9). The effect of the generalisability vs verifiability trade-off can thus be rather severe for geometric subspaces.
This experiment demonstrates the importance of using the generalisability metric: if one only took into account the verifiability of the subspaces one would choose $\mathbb{H}_{\epsilon=0.005}$, obtaining mathematically sound but pragmatically useless results. We argue that this is a strong argument for including generalisability as a standard metric in reporting NLP verification results in the future.
4.4 Verifiability-Generalisability Trade-off for Semantic Subspaces
The previous subsection has shown that the verifiability-generalisability trade-off is not resolvable by geometric manipulations alone. In this section we argue that using semantic subspaces can help to improve the effects of the trade-off. The main hypothesis that we are testing is: semantic subspaces constructed using semantic-preserving perturbations are more precise, and this in turn improves both verifiability and generalisability.
We will use the construction given in Definition 6. As Table 11 illustrates, we construct several semantic hyper-rectangles on sentences of the positive class using character-level ($\mathbb{H}_{char}$, $\mathbb{H}_{del.}$, $\mathbb{H}_{ins.}$, $\mathbb{H}_{rep.}$, $\mathbb{H}_{repl.}$, $\mathbb{H}_{swap.}$), word-level ($\mathbb{H}_{word}$) and sentence-level perturbations ($\mathbb{H}_{pj}$). The subscripts char and word refer to the kind of perturbation algorithm, while del., ins., rep., repl., swap. and pj refer to the type of perturbation, where pj stands for Polyjuice (see Section 3.1). Notice comparable volumes of all these shapes, and compare with $\mathbb{H}_{\epsilon=0.05}$.
Experiment name Hyper-rectangles construction method Avg. volume of hyper-rectangles RUAR Number of hyper-rectangles RUAR Avg. volume of hyper-rectangles Medical Number of hyper-rectangles Medical
$\mathbb{H}_{char}$ Set of hyper-rectangles for character perturbations 1.54e-30 3400 7.66e-31 989
$\mathbb{H}_{word}$ Set of hyper-rectangles for word perturbations 1.28e-30 3400 - -
$\mathbb{H}_{pj}$ Set of hyper-rectangles for polyjuice sentence perturbations - - 2.01e-28 989
$\mathbb{H}_{swap.}$ Set of hyper-rectangles for swapping perturbations 1.57e-31 3400 3.42e-31 989
$\mathbb{H}_{repl.}$ Set of hyper-rectangles for replacement perturbations 9.84e-31 3400 3.43e-31 989
$\mathbb{H}_{del.}$ Set of hyper-rectangles for deletion perturbations 3.46e-31 3400 1.24e-32 989
$\mathbb{H}_{ins.}$ Set of hyper-rectangles for insertion perturbations 3.21e-31 3400 9.11e-33 989
$\mathbb{H}_{rep.}$ Set of hyper-rectangles for repetition perturbations 1.56e-31 3400 1.06e-32 989
4.4.1 Verifiability of Semantic Subspaces
We pass each set of hyper-rectangles and the network $N_{base}$ to the verifiers ERAN and Marabou to measure verifiability of the subspaces. Table 12 illustrates the verification results obtained using ERAN. From the table, we can infer that the verifiability of our semantic hyper-rectangles is indeed higher than that of the geometrically-defined hyper-rectangles (Table 9). Furthermore, our semantic hyper-rectangles, while unable to reach the verifiability of $\mathbb{H}_{\epsilon=0.005}$, achieve notable higher verification than its counterpart of comparable volume $\mathbb{H}_{\epsilon=0.05}$. From this experiment, we conclude that not only volume, but also precision of the subspaces has an impact on their verifiability.
Dataset Model $\boldsymbol{\mathbb{H}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}_{word}}$ $\boldsymbol{\mathbb{H}_{char}}$ $\boldsymbol{\mathbb{H}_{del.}}$ $\boldsymbol{\mathbb{H}_{ins.}}$ $\boldsymbol{\mathbb{H}_{rep.}}$ $\boldsymbol{\mathbb{H}_{repl.}}$ $\boldsymbol{\mathbb{H}_{swap.}}$ $\boldsymbol{\mathbb{H}_{pj}}$
RUAR $N_{base}$ 0.00% 1.80% 0.87% 1.62% 2.63% 1.66% 0.94% 2.07% -
Medical $N_{base}$ 5.00% - 39.71% 39.62% 44.66% 48.71% 37.49% 42.60% 50.09%
Table 12: Verifiability of the baseline DNN on the RUAR and the Medical datasets, for a selection of semantic subspaces; using the ERAN verifier.
Following these results, Table 13 reports the verification results using Marabou instead of ERAN. As shown, Marabou is able to verify up to $66.83\%$ ($\mathbb{H}_{rep.}$), while ERAN achieves at most $50.09\%$. This shows that Marabou outperforms ERAN. This is most likely due to the fact that Reluplex algorithm of Marabou achieves better precision on ReLU networks, that we use. Overall, the Marabou experiment confirms the trends of improved verifiability shown by ERAN and thus confirms our hypothesis about importance of shape precision.
Dataset Model $\boldsymbol{\mathbb{H}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}_{word}}$ $\boldsymbol{\mathbb{H}_{char}}$ $\boldsymbol{\mathbb{H}_{del.}}$ $\boldsymbol{\mathbb{H}_{ins.}}$ $\boldsymbol{\mathbb{H}_{rep.}}$ $\boldsymbol{\mathbb{H}_{repl.}}$ $\boldsymbol{\mathbb{H}_{swap.}}$ $\boldsymbol{\mathbb{H}_{pj}}$
RUAR $N_{base}$ 1.79% 11.69% 4.88% 4.35% 9.72% 9.46% 5.65% 8.07% -
Medical $N_{base}$ 37.96% - 64.03% 64.15% 64.65% 66.83% 64.75% 64.36% 61.57%
Table 13: Verifiability of the baseline DNN on the RUAR and the Medical datasets, for a selection of semantic subspaces; using the Marabou verifier.
4.4.2 Generalisability of Semantic Subspaces
It remains to establish whether the more verifiable semantic subspaces are also more generalisable. Whereas Table 10 compared the generalisability of $\mathbb{H}_{\epsilon=0.005}$ and $\mathbb{H}_{\epsilon=0.05}$ with that of $\mathbb{H}_{sh}$, Table 14 compares their generalisability to the most verifiable semantic subspaces, $\mathbb{H}_{word}$ and $\mathbb{H}_{pj}$. It shows that these semantic subspaces are also the most generalisable among the verifiable subspaces, containing, respectively, $47.67\%$ and $28.74\%$ of the unseen sentences. Note that among all the experiments, only $\mathbb{H}_{sh}$ has higher generalisability, but its verifiability is 0.
Dataset Experiment Avg. Volume of hyper-rectangles Generalisability (%) Number of sentences contributing to generalisability Total Sentences in $\mathcal{A}^{b}_{t}(\mathcal{Y}^{pos})$
RUAR $\mathbb{H}_{\epsilon=0.005}$ 1.00e-60 1.95 2821 144500
$\mathbb{H}_{\epsilon=0.05}$ 1.00e-30 38.47 55592 144500
$\mathbb{H}_{word}$ 1.28e-30 47.67 68882 144500
Medical $\mathbb{H}_{\epsilon=0.005}$ 1.00e-60 0.09 10 11209
$\mathbb{H}_{\epsilon=0.05}$ 1.00e-30 28.49 3194 11209
$\mathbb{H}_{pj}$ 2.01e-28 28.74 3222 11209
Table 14: Generalisability of the selected geometric subspaces $\mathbb{H}_{\epsilon=0.005}$ and $\mathbb{H}_{\epsilon=0.05}$ and the semantic subspaces $\mathbb{H}_{word}$ and $\mathbb{H}_{pj}$, measured on the sets of semantic perturbations $\mathcal{A}^{b}_{t_{RUAR}}(\mathcal{Y}^{pos})$ and $\mathcal{A}^{b}_{t_{medical}}(\mathcal{Y}^{pos})$. Note that the generalisability of $\mathbb{H}_{sh}$ (Table 10), despite it having the volume 19 order of magnitudes bigger, is only $3\%$ greater than $\mathbb{H}_{word}$.
@@ 4.4.2 Generalisability of Semantic Subspaces
Dataset Experiment Total Volume of Hyper-rectangles Training Embedding Space Covered (%)
RUAR $\mathbb{H}_{\epsilon=0.005}$ 2.89e-57 4.71e-53
$\mathbb{H}_{\epsilon=0.05}$ 2.89e-27 4.71e-23
$\mathbb{H}_{word}$ 3.7e-27 6.03e-23
Medical $\mathbb{H}_{\epsilon=0.005}$ 9.89e-58 6.92e-53
$\mathbb{H}_{\epsilon=0.05}$ 9.89e-28 6.92e-23
$\mathbb{H}_{pj}$ 1.63e-25 1.14e-20
Table 15: Total volume and percentage of the training embedding space covered by our best hyper-rectangles. The total volume of the training embedding space for RUAR is $56.14e-$ 5, and for Medical is $51.43e-$ 5.
4.5 Adversarial Training on Semantic Subspaces
In this section, we study the effects that adversarial training methods have on the verifiability of the previously defined subspaces in Tables 8 and 11. By comparing the effectiveness of the different training approaches described in Section 3.5, we show in this section that adversarial training based on our new semantic subspaces is the most efficient. Three kinds of training are deployed in this section:
  1. 1.
    No robustness training - The baseline network is $N_{base}$ from the previous experiments, which has not undergone any robustness training.
  2. 2.
    Data augmentation. We obtain three augmented datasets $\mathcal{Y}\cup\mathcal{A}^{5}_{char}(\mathcal{Y}^{pos})$, $\mathcal{Y}\cup\mathcal{A}^{6}_{word}(\mathcal{Y}^{pos})$ and $\mathcal{A}^{5}_{pj}(\mathcal{Y}^{pos})$ where $\mathcal{A}(\cdot)$ is defined in Section 4.4. The subscripts char and word denote the type of perturbation as detailed in Tables 4 and 5, while the subscript pj refers to the sentence level perturbations generated with Polyjuice. We train the baseline architecture, using the standard stochastic gradient descent and cross entropy loss, on the augmented datasets, and obtain DNNs $N_{char-aug}$, $N_{word-aug}$ and $N_{pj-aug}$.
  3. 3.